# Foldings in relatively hyperbolic groups

### v1

@@ 4.1 Equivalence classes of carrier graphs of groups
    If all peripheral subgroups are Abelian then $\alpha_{e}(A_{e})$ is a subgroup of $A_{c_{i}}$.
@@
If $\mathbb{A}$ is minimal, Lemma 2.25 gives an upper bound for the number $N(\mathbb{A})$ of maximal trivial segments of $\mathbb{A}$, which only depends on the equivalence class of $\mathcal{A}$.
@@
Let $(G,\mathbb{P})$ be a torsion-free toral relatively hyperbolic groups. For any $(G,\mathbb{P})$-carrier graph of groups $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ and $i\in\{1,\ldots,k\}$ denote by $\bar{A}_{c_{i}}$ the root closure of
@@
(1) Clearly $\eta_{i}|_{\tilde{A}_{c_{i}}}=\hbox{id}|_{\tilde{A}_{c_{i}}}$. The claim follows as $\eta_{i}|_{\bar{A}_{c_{i}}}$ is uniquely determined by $\eta_{i}|_{\tilde{A}_{c_{i}}}$ by Lemma 4.5.
@@ 4.2 Folds and AO-Moves
If the conclusion of Corollary 3.6 does not hold, i.e. if at least one of the cases (A1)-(A8) occurs, then the carrier graph can be altered in some controlled way discussed below. In cases (A1)-(A5) and (A7) these modification can be thought of (up to some preprocessing) as folds of graphs of groups as introduced by Bestvina and Feighn [8] and Dunwoody [20], see also [29]. In case (A6) a variation of a move introduced by Arzhantseva and Olshanskii in the context of graphs can be applied, [6]. In case (A8) the modification is merely cosmetic.
The alterations are designed in such a way that they simplify the carrier graph of groups by decreasing the values of certain invariant. Among them is the following complexity associated to a $(G;\mathbb{P})$-carrier graph $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$:
$c(\mathcal{A}):=\sum_{e\in E_{0},\ e\ non-peripheral}|g_{e}|_{X\cup\mathcal{P}}+\sum_{1\leq i\leq k}\hbox{val}(c_{i})$
In the remainder of this Section folds and other moves are introduced that modify a $(G,\mathbb{P})$-carrier graph $\mathcal{A}$ if one of the situations (A1)-(A8) spelled out in Corollary 3.6 occurs. These modifications preserve the subgroup of $G$ that is represented (at least up to conjugacy as we don’t distinguish base points). The following is always assumed without explicit mentioning:
@@
We may further assume that $L\geq\nu(G,\mathbb{P},X,M)$ and therefore $L>|g_{e}|_{X}$ for all essential edges $e\in EA$. An analysis of the proof of Corollary 3.6 actually implies that this also follows from the choice of the various constants.
@@
      If $i=j$: Delete $s$ and replace $A_{c_{i}}$ by $\langle A_{c_{i}},g\rangle$. The free rank decreases by one.
@@
    Suppose $t$ is such a path of minimal length. If $t$ has an edge with trivial edge group, let $e$ be such an edge. Otherwise let $e$ be a non-peripheral edge of $t$.
@@
    • If $e$ is essential then the number of essential edges in the core of the only affected free factor decreases by at least one.
@@
    • Either the free rank or the number of free factors decreases by one. The new $(G,\mathbb{P})$-carrier graph is $\max(3L,M)$-prenormal. This happens in (b).
@@
    Replace $\mathcal{A}$ by $\mathcal{A}^{\prime}$ in the following way: Delete $t^{\prime}$ and add a new free edge $e$ with $\alpha(e)=t_{-}$, $\omega(e)=t_{+}$, $g_{e}=\nu_{\mathcal{A}}(\hat{t})$ and $A_{e}=1$. This strictly decreases the complexity of $\mathcal{A}$.
    Upshot: Free rank and all free factors are preserved, thus the new carrier graph is again $M$-prenormal and strongly equivalent to $\mathcal{A}$. Moreover $c(\mathcal{A}^{\prime})<c(\mathcal{A})$.
@@
    Choose a geodesic word $upv$ with $p\in P_{i}$ for some $i$. Subdivide $e$ twice to obtain a segment consisting of the edges $e_{1}$, $e_{2}$, $e_{3}$ and $e_{4}$ with labels $u$, $p$, $1$ and $v$ respectively. Declare the subcarrier graph consisting of $e_{2}$ and $e_{3}$ to be an additional peripheral star with central vertex $\omega(e_{2})=\alpha(e_{3})$.
@@
    The free rank decreases.
@@
    The free rank is preserved and the number of free factors decreases.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has finite symmetric generating set $X$. Let $\mathfrak{C}$ be an equivalence class of $M$-prenormal $(G,\mathbb{P})$-carrier graphs.
@@ 4.3 Finiteness Theorems
Let $G$ be a finitely generated torsion-free group that is hyperbolic relative to a family =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of noetherian groups. Let $n\in\mathbb{N}$. Then there exist finitely many equivalence classes $\mathfrak{C}_{1},\ldots,\mathfrak{C}_{l}$ of $(G,\mathbb{P})$-carrier graphs such that for any locally relatively quasiconvex, $n$-generated subgroup $U$ of $G$ there exists some $j\in\{1,\ldots,l\}$ and $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})\in\mathfrak{C}_{j}$ such that the following hold:
@@
One now continues this argument for each of these finitely many equivalence classes. In this process case (1) and case (2) of Lemma 4.7 can clearly only occur finitely many times as new free factors can only emerge if the free rank decreases. In the remaining case the edge group of at least one essential edge of the core increases (and some possibly disappear). As the $P_{i}$ are assumed to be noetherian, this can also only happen finitely many times. The first claim now follows from König’s Lemma.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to a family =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of noetherian groups. Let $n\in\mathbb{N}$.

### v2

@@ 4.1 Equivalence classes of carrier graphs of groups
    If in addition all peripheral subgroups are Abelian then $\alpha_{e}(A_{e})$ is a subgroup of $A_{\alpha(e)}=A_{c_{i}}$.
@@
Lemma 2.29 gives an upper bound for the number $N(\mathcal{A})$ of maximal trivial segments of $\mathbb{A}$, which only depends on the equivalence class of $\mathcal{A}$.
@@
Let $(G,\mathbb{P})$ be a torsion-free toral relatively hyperbolic group. For any $(G,\mathbb{P})$-carrier graph of groups $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ and $i\in\{1,\ldots,k\}$ denote by $\bar{A}_{c_{i}}$ the root closure of
@@
(1) For any essential edge $e$ with $\alpha(e)\in C_{i}$ the definition of equivalence implies that $\eta_{i}(\alpha_{e}(A_{e}))\subset P_{m_{i}}$ is in $G$ conjugate to $\alpha_{f(e)}(B_{f(e)})\subset P_{m_{i}}$. As $P_{m_{i}}$ is Abelian and malnormal in $G$ it follows that $\eta_{i}(\alpha_{e}(A_{e}))=\alpha_{f(e)}(B_{f(e)})$. Thus $\tilde{A}_{c_{i}}=\tilde{B}_{c_{i}^{\prime}}$ and therefore $\bar{A}_{c_{i}}=\bar{B}_{c_{i}^{\prime}}$ by Lemma 4.5.
Clearly $\eta_{i}|_{\tilde{A}_{c_{i}}}=\hbox{id}|_{\tilde{A}_{c_{i}}}$. The last claim now follows as $\eta_{i}|_{\bar{A}_{c_{i}}}$ is uniquely determined by $\eta_{i}|_{\tilde{A}_{c_{i}}}$ by Lemma 4.5.
@@ 4.2 Folds and AO-Moves
If the conclusion of Corollary 3.6 does not hold, i.e. if at least one of the cases (A1)-(A8) occurs, then the carrier graph can be altered in some controlled way as discussed below. In cases (A1)-(A5) and (A7) these modification can be thought of (up to some preprocessing) as folds of graphs of groups as introduced by Bestvina and Feighn [9] and Dunwoody [22], see also [31]. In case (A6) a variation of a move introduced by Arzhantseva and Olshanskii in the context of graphs can be applied, see [6]. In case (A8) the modification is merely cosmetic.
These alterations are designed in such a way that they simplify the carrier graph of groups by decreasing the values of certain invariants. Among them is the following complexity associated to a $(G,\mathbb{P})$-carrier graph $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$:
$c(\mathcal{A}):=\sum_{\begin{subarray}{c}e\in E_{0}\\ \ e\ non-peripheral\end{subarray}}|g_{e}|_{X\cup\mathcal{P}}+\sum_{1\leq i\leq k}\hbox{val}(c_{i})$
In the remainder of this Section folds and other moves are introduced that modify a $(G,\mathbb{P})$-carrier graph $\mathcal{A}$ if one of the situations (A1)-(A8) spelled out in Corollary 3.6 occurs. These modifications preserve the subgroup of $G$ that is represented, i.e. the image of the map $\nu_{\mathcal{A}}:\pi_{1}(\mathbb{A},v_{0})\to G$. The following is always assumed without explicit mentioning:
@@
Let $L$ be the constant from the conclusion of Corollary 3.6. We may assume that $L\geq\nu(G,\mathbb{P},X,M)$ and therefore $L>|g_{e}|_{X}$ for all essential edges $e\in EA$. An analysis of the proof of Corollary 3.6 actually implies that this also follows from the choices of the various constants.
@@
      If $i=j$: Delete $s$ and replace $A_{c_{i}}$ by $\langle A_{c_{i}},g\rangle$.
@@
    Suppose $t$ is such a path of minimal length. If $t$ has an edge with trivial edge group, let $e$ be such an edge. Otherwise $t$ is a path in some free factor of $\mathbb{A}$ whose initial and terminal edges are essential connecting $v$ and $v^{\prime}$ to the same non-trivial peripheral star. Let $e$ be initial essential edge of $t$.
@@
    • If $e$ is essential then the number of essential edges in the core of the only affected free factor decreases by at least one. Note that the prenormalization process might add new essential edges, however these added essential edges will not lie in the core of affected free factor.
@@
    • Either the relative rank or the number of free factors decreases by one. This happens in (b).
@@
[…]
    Upshot: The relative rank and all free factors are preserved, thus the new carrier graph is again $M$-prenormal and strongly equivalent to $\mathcal{A}$. Moreover $c(\mathcal{A}^{\prime})<c(\mathcal{A})$.
@@
    Choose a geodesic word upvupv with $p\in P_{i}$ for some $i$. Subdivide $e$ twice to obtain a segment consisting of the edges $e_{1}$, $e_{2}$, $e_{3}$ and $e_{4}$ with labels $u$, pp, 11 and $v$ respectively. Declare the subcarrier graph consisting of $e_{2}$ and $e_{3}$ to be an additional peripheral star with central vertex $\omega(e_{2})=\alpha(e_{3})$.
@@
    The relative rank decreases.
@@
    The relative rank is preserved and the number of free factors decreases.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has a finite symmetric generating set $X$. Let $\mathfrak{C}$ be an equivalence class of $M$-prenormal $(G,\mathbb{P})$-carrier graphs.
@@ 4.3 Finiteness Theorems
Let $G$ be a finitely generated torsion-free group that is hyperbolic relative to a family $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of noetherian groups. Let $n\in\mathbb{N}$. Then there exist finitely many equivalence classes $\mathfrak{C}_{1},\ldots,\mathfrak{C}_{l}$ of $(G,\mathbb{P})$-carrier graphs such that for any locally relatively quasiconvex, $n$-generated subgroup $U$ of $G$ there exists some $j\in\{1,\ldots,l\}$ and $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})\in\mathfrak{C}_{j}$ such that the following hold:
@@
One now continues this argument for each of these finitely many equivalence classes. In this process case (1) and case (2) of Lemma 4.7 can clearly only occur finitely many times as new free factors can only emerge if the relative rank decreases. In the remaining case the edge group of at least one essential edge of the core increases (and some possibly disappear). As the $P_{i}$ are assumed to be noetherian, this can also only happen finitely many times. The first claim now follows from König’s Lemma.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to a family $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of noetherian groups. Let $n\in\mathbb{N}$.